# The Physics of Local Optimization in Complex Disordered Systems

### v1

@@ Abstract
To systematically study the relation of local and global solutions in spin systems, we rely on exact (combinatorial) GS algorithms, using a minimum-weight perfect matching (MWPM) method [29] based on Blossom V [30] for the polynomial 2D problem and the general-purpose Gurobi software [31] for 3D systems and the SK model. Figure 1 schematically depicts the setup: we computed GSs for the entire ($N$-spin) systems as well as for subsystems with $N_{\mathrm{sub}}$ lattice sites centered on the considered bond $\langle i,j\rangle$ (corresponding to the LSBS), comparing the predicted values of the product $\sigma_{i_{0}}\sigma_{j_{0}}$ (open boundary conditions are applied in both cases). If the results agree, the bond product was predicted correctly (with zero error). For $L\times L$ square lattices of sizes $L=256$, $512$, $1024$, we examined $L_{\rm sub}\times L_{\rm sub}$ subsystems with $4\leq L_{\rm sub}\leq 512$, using $\sim 60,000$ disorder realizations. For cubic lattices of side lengths $L=10$ and $12$, the $L_{\rm sub}\times L_{\rm sub}\times L_{\rm sub}$ subsystems were of size $4\leq L_{\rm sub}\leq 10$, and we used $\sim 2,000$ realizations. From these calculations we estimate the LSBS error rate $\mathcal{E}$. Figure 2 shows the dependence of $\mathcal{E}$ on the subsystem size for several cases. For the $\overline{J}=0$ lattice systems we find that, although the 2D and 3D spin-glass GS problems belong to different complexity classes (P and NP, respectively), in both cases the error rate is well described by a power-law decay in the subsystem size  [^1] While we will largely focus on the central nearest neighbor bond (with $i=i_{0}$ and $j=j_{0}$), in the End Matter we discuss more general situations in which the sites $i$ and $j$ are a distance ${\sf r}_{ij}>1$ apart. To that end and in order to avoid cumbersome notation, we will often dispense with “0” subscripts.  [^2] In 2D, for the largest values of $L_{\rm sub}={\cal{O}}(L)$, the righthand side of Eq. (2) constitutes an upper bound on the error $\mathcal{E}_{ij}$. Clearly, when $L_{\rm sub}=L$ (when the subsystem is the entire system), the error $\mathcal{E}_{ij}=0$ by definition., viz.
@@
Figure 3 illustrates the sensitivity $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ of each bond of a given $L=128$ square-lattice sample, and how it relates to the error rate of the LSBS. As is seen here (and in Fig. 4, End Matter), the LSBS error rate decays with increasing $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$. We find this decay to be described well by a stretched-exponential form
$\mathcal{E}_{ij}\sim k_{J}\,e^{-\alpha|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|^{\beta}}.$ (3)
The parameters for $L=1024$, $L_{\rm sub}=8$ square and $L=12$, $L_{\rm sub}=4$ cubic lattice systems appear in Table 1. The modulus $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ is equal to twice the excitation energy $\Delta E$, cf. Eq. (LABEL:eq:jccompute) in Ref. [28], Sec. LABEL:sec:compute_jc. Thus the accuracy of the LSBS for a certain bond is determined by the respective excitation energy of the bond in this subsystem.
We further tuned the central bonds of $L=128$ square and $L=12$ cubic lattices to their critical values and then computed the respective critical threshold changes ($\Delta J_{\mathrm{c},ij}$) of all other bonds. We find that the average difference $|\Delta J_{\mathrm{c},ij}|$ decays algebraically with distance from the modified central bond (see Fig. 5, End Matter),

### v2

@@ Abstract
To systematically study the relation of local and global solutions in spin systems, we rely on exact (combinatorial) GS algorithms, using a minimum-weight perfect matching (MWPM) method [29] based on Blossom V [30] for the polynomial 2D problem and the general-purpose Gurobi software [31] for 3D systems and the SK model. Fig.  1 schematically depicts the setup: we computed GSs for the entire ($N$-spin) systems as well as for subsystems with $N_{\mathrm{sub}}$ lattice sites centered on the considered bond $\langle i,j\rangle$ (corresponding to the LSBS), comparing the predicted values of the product $\sigma_{i_{0}}\sigma_{j_{0}}$ (open boundary conditions are applied in both cases). If the results agree, the bond product was predicted correctly (with zero error). For $L\times L$ square lattices of sizes $L=256$, $512$, $1024$, we examined $L_{\rm sub}\times L_{\rm sub}$ subsystems with $4\leq L_{\rm sub}\leq 512$, using $\sim 60,000$ disorder realizations. For cubic lattices of side lengths $L=10$ and $12$, the $L_{\rm sub}\times L_{\rm sub}\times L_{\rm sub}$ subsystems were of size $4\leq L_{\rm sub}\leq 10$, and we used $\sim 2,000$ realizations. From these calculations we estimate the LSBS error rate $\mathcal{E}$. Fig. 2 shows the dependence of $\mathcal{E}$ on the subsystem size for several cases. For the $\overline{J}=0$ lattice systems we find that, although the 2D and 3D spin-glass GS problems belong to different complexity classes (P and NP, respectively), in both cases the error rate is well described by a power-law decay in the subsystem size  [^1] While we will largely focus on the central nearest neighbor bond (with $i=i_{0}$ and $j=j_{0}$), in the End Matter we discuss more general situations in which the sites $i$ and $j$ are a distance ${\sf r}_{ij}>1$ apart. To that end and in order to avoid cumbersome notation, we will often dispense with “0” subscripts.  [^2] In 2D, for the largest values of $L_{\rm sub}={\cal{O}}(L)$, the righthand side of Eq. (2) constitutes an upper bound on the error $\mathcal{E}_{ij}$. Clearly, when $L_{\rm sub}=L$ (when the subsystem is the entire system), the error $\mathcal{E}_{ij}=0$ by definition., viz.
@@
Fig. 3 illustrates the sensitivity $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ of each bond of a given $L=128$ square-lattice sample, and how it relates to the error rate of the LSBS. As seen here (and in Fig. 4, End Matter), the LSBS error rate decays with increasing $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$. We find this decay to be described well by, respectively, the compressed-exponential/exponential forms
$\mathcal{E}_{ij}\sim k_{J}\,e^{-(\frac{|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|}{\sf{J}})^{\beta}},$ (3)
with $\beta>1$ (2D) and $\beta=1$ (3D), see Table 1. As in Eq. (2), in both 2D and 3D EA systems, $\mathcal{E}_{ij}$ exhibits a similar functional form dependence on the deviation of the coupling of the investigated bond from its critical value (Ref. [28], Sec. LABEL:sec:2Dvs3D). This latter modulus $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ is equal to twice the excitation energy $\Delta E$, cf. Eq. (LABEL:eq:jccompute) in Ref. [28], Sec. LABEL:sec:compute_jc. Thus the accuracy of the LSBS for a certain bond is determined by the respective excitation energy of the bond in this subsystem.
We further tuned the central bonds of the $L=128$ square and $L=12$ cubic lattices to their critical values and then calculated the respective changes in critical threshold ($\Delta J_{\mathrm{c},ij}$) of all other bonds. We find that the average difference $|\Delta J_{\mathrm{c},ij}|$ decays algebraically with distance from the modified central bond (see Fig. 5, End Matter),